OpenStudy (anonymous):

9th term in the geometric sequence 4, 12, 36, ...

OpenStudy (anonymous):

\[X_{9}=ar ^{n-1}\] a=the first term r=the common ratio between terms \[X_{9}=4(3)^{9-1}\]\[X_{9}=12^{8}\]\[X_{9}=429981696\]